A pension plan is obligated to make disbursements of $2.0 million, $3.0 million, and $2.0 million at the end of each of the next three years, respectively. Find the duration of the plan's obligations if the interest rate is 11% annually. *the solution gives a table but I'm unsure how to figure out pv. explanation on financial calculator would be helpful.
Answer #1

Rate =11%
Using Financial Calculator
I/Y=11%,CF1 =2;CF2=3;CF3=2, CPT PV =5.6991 or 5.70
Using Formula
PV =2/1.11+3/1.11^2+2/1.11^3 =5.6691 or 5.70

Duration =(1*1/1.11+3*2/1.11^2+2*3/1.1^3)/5.6991 =1.80 years